# Schrödinger as a Quantum Programmer: Estimating Entanglement via Steering

### v3

@@ 1 Introduction
where $\{|x\rangle_{R}\}_{x\in\mathcal{X}}$ is an orthonormal basis. Purifications are not unique, but all other purifications of $\sigma_{AB}$ are related to the one in (2) by the action of a unitary operation on the reference system $R$ [40]. By inspecting (2), we see that the systems $A$ and $B$ can be steered into the probabilistic ensemble $\{(p(x),\psi^{x}_{A}\otimes\phi^{x}_{B})\}_{x\in\mathcal{X}}$ of product states by performing the projective measurement $\{|x\rangle\!\langle x|_{R}\}_{x\in\mathcal{X}}$ on the reference system $R$ of $\varphi_{RAB}$. This leads to an idea for testing separability in the general case. If purification of a general state $\rho_{AB}$ is available and the state $\rho_{AB}$ is indeed separable, then one can a) try to find the unitary that realizes the purification in (2) and b) perform the measurement $\{|x\rangle\!\langle x|_{R}\}_{x\in\mathcal{X}}$ on the reference system $R$. After receiving the outcome $x$, one can finally test whether the reduced state is a product state.
@@
In this paper, we expand on the abovementioned idea to develop various separability tests using the quantum steering effect. Our separability test for mixed states consists of a distributed quantum computation involving two parties: a computationally unbounded server, called a prover, which can, in principle, perform any quantum computation imaginable, and a computationally limited client, called a verifier, which can perform time-efficient quantum computations (see Figure 1). We show in Theorems 1 and 2 that the acceptance probabilities of our algorithms, in the ideal case, are directly related to a bonafide entanglement measure, the fidelity of separability. We also employ concepts from quantum computational complexity theory [41, 42] to understand how difficult this test is to perform. Our second contribution results from a modification of our separability test. In an attempt to design a practical algorithm, we replace the prover with a combination of parameterized unitary circuits and classical optimization techniques to perform the necessary computation. This results in a variational quantum steering algorithm (VQSA) that approximates the aforementioned separability test (see Figure 2). The concept of quantum steering is again at the heart of our VQSA, just like the test for separability that it approximates. Interestingly, we prove that the acceptance probability of both tests is related to an entanglement measure called fidelity of separability [22, 23]. We also generalize our separability test and VQSA to the multipartite setting using appropriate definitions of multipartite separability.
@@
Our approach is distinct from recent work on quantum algorithms for estimating entanglement. For example, VQAs have been used to address this problem by estimating the Hilbert–Schmidt distance [45], by creating a zero-sum game using parameterized unitary circuits [46], by employing symmetric extendibility tests [31], by estimating logarithmic negativity [47], and using the positive map criterion [47]. VQAs have also been used to estimate the geometric measure of entanglement of multiqubit pure states [48]. The work of [49] is the closest related to ours, but the test used there requires two copies of the state of interest and controlled swap operations, while our VQSA does not require either. In contrast, we introduce a paradigm for VQAs involving parameterized mid-circuit measurements, which is the core of our method for estimating entanglement, and we suspect that this approach will be helpful in future work for a wide variety of VQAs. Furthermore, as we show in Theorems 1 and 2, the acceptance probabilities of our algorithms, in the ideal case, are directly related to a bonafide entanglement measure, the fidelity of separability.
@@ 2.1 Quantum Interactive Proof for Fidelity of Separability
where $\Pi_{A^{\prime}A}^{\operatorname{sym}}$ is the projector onto the symmetric subspace of the $A^{\prime}$ and $A$ systems, and $\operatorname{EB}_{R\to A^{\prime}}$ denotes the set of all entanglement-breaking channels with input system $R$ and output system $A^{\prime}$. We find in Theorem 1 below that the maximum acceptance probability in (4) can be expressed as a simple function of the fidelity of separability of $\rho_{AB}$, the latter defined as [22, 23]
@@
For a pure state $\psi_{RAB}$, the following holds
@@ 2.2 Variational Quantum Steering Algorithm for Fidelity of Separability
Our VQSA is structured as follows. Let $\rho_{AB}$ denote the state for which we want to estimate the fidelity of separability, and let $\psi_{RAB}$ be a purification of it, which results from the action of the unitary operator $U^{\rho}$ on the all-zeros pure state $|0\rangle\!\langle 0|$. Once we have $\psi_{RAB}$, we can attempt to access all possible pure-state decompositions $\{(p(x),\psi_{AB}^{x})\}_{x\in\mathcal{X}}$ of $\rho_{AB}$ by acting on system $R$ with unitary operations. We use the first parameterized unitary $W_{R}(\Theta)$ to do so. To ensure that we have a sufficient number of measurement outcomes (to cover the possible case when $|\mathcal{X}|=\text{rank}(\rho_{AB})^{2}$), we can prepare some ancilla qubits in the all-zeros state, for a system $R^{\prime}$, and act with $W$ on $R$ and $R^{\prime}$. However, without loss of generality, these extra qubits can be grouped as part of an overall reference system, relabeled as $R$.
@@ 2.3 Benchmarking Semidefinite Programs and Examples
As a second example, we consider a state consisting of four qubits. Let us consider the four qubit state $|\psi\rangle$ defined as follows:
@@ 2.4 Generalization to Multipartite Fidelity of Separability
We also generalize our VQSA to measure the fidelity of separability of multipartite states in the following fashion. A multipartite state $\rho_{A_{1}\cdots A_{M}}\in\mathcal{D}(\mathcal{H}_{A_{1}\cdots A_{M}})\equiv\mathcal{D}(\mathcal{H}_{A_{1}}\otimes\cdots\otimes\mathcal{H}_{A_{M}})$ is separable if it can be written as
@@
where $\{\mu^{x}_{R}\}_{x}$ is a rank-one POVM and $\{\phi^{x,i}_{A^{\prime}_{i}}\}_{x,i}$ is a set of pure states. The prover sends systems $(A^{M-1})^{\prime}\equiv A^{\prime}_{1}\cdots A^{\prime}_{M-1}$ to the verifier. Finally, the verifier performs a collective swap test on these systems and the systems $A_{1}\cdots A_{M}$, as depicted in Figure 5. The acceptance probability of this distributed quantum computation is given by
@@
where $\Pi_{(A^{M-1})^{\prime}A^{M-1}}^{\operatorname{sym}}$ is the projection onto the symmetric subspace of systems $(A^{M-1})^{\prime}$ and $A^{M-1}$ and $\operatorname{EB}_{M-1}$ denotes the set of multipartite entanglement-breaking channels defined in (11). This leads to the following theorem:
@@
Figure 6: VQSA to estimate the multipartite fidelity of separability $F_{s}(\rho_{A_{1}A_{2}A_{3}A_{4}})$. The unitary circuit $U^{\rho}$ prepares the state $\psi_{RA_{1}A_{2}A_{3}A_{4}}$, which is a purification of $\rho_{A_{1}A_{2}A_{3}A_{4}}$. The parameterized circuit $W_{R}(\Theta)$ acts on $R$ to evolve the state to another purification of $\rho_{A_{1}A_{2}A_{3}A_{4}}$. The following measurement, labeled “steering measurement,” steers the remaining systems to be in a state $\psi_{A_{1}A_{2}A_{3}A_{4}}^{x}$ if the measurement outcome $x$ occurs. Conditioned on the outcome $x$, the final parameterized circuits $U^{x,1}_{A_{1}}(\Theta^{x}_{1})$, $U^{x,2}_{A_{2}}(\Theta^{x}_{2})$, and $U^{x,3}_{A_{3}}(\Theta^{x}_{3})$ are applied and the subsequent measurement accepts with a maximum probability of $F_{s}(\rho_{A_{1}A_{2}A_{3}A_{4}})$.
@@ 2.5 Quantum Computational Complexity Considerations
$\max_{\begin{subarray}{c}\{(p(x),\psi^{x})\}_{x},\left\{\varphi^{x}\right\}_{x},\\ \rho_{B}=\sum_{x}p(x)\psi_{B}^{x}\end{subarray}}\sum_{x}p(x)F(\psi_{B}^{x},\mathcal{N}_{A\rightarrow B}(\varphi_{A}^{x})).$ (15)
See Appendix J for details and an interpretation of this quantity.
@@ 3 Conclusion and Discussion
In this paper, we detailed a distributed quantum computation to test the separability of a quantum state that, at its core, uses quantum steering. This test demonstrated a link between quantum steering and the separability problem. The acceptance probability of this distributed quantum computation is directly related to an entanglement measure known as the fidelity of separability. Using the test’s structure, we also showed computational complexity-theoretic results and established a link between quantum steering, quantum algorithms, and quantum computational complexity. By replacing the prover with a parameterized circuit, we modified this distributed quantum computation to develop our VQSA, a novel kind of variational quantum algorithm that uses quantum steering to address the problem of estimating the fidelity of separability. This algorithm allows for the direct estimation of the fidelity of separability without the need for state tomography and subsequent approximate tests on separability. Our algorithm is not unitary due to the mid-circuit measurement on system $R$ and the consequent conditional operation applied on system $A$. This is an important distinction from most VQAs, which do not use a parameterized mid-circuit measurement. We also discuss multipartite generalizations of both our separability test and VQSA. Finally, we simulated our VQSA using the noisy Qiskit Aer simulator, which showed favorable convergence trends and was compared against two classical SDP benchmarks.
@@
Our VQSA has applications beyond entanglement quantification on a single quantum computer. We can also think of our VQSA as a distributed variational quantum algorithm for measuring the entanglement of a bipartite state. See [58, 59, 60] for previous instances of distributed VQAs. Indeed, our algorithm can be executed over a quantum network in which each node has quantum and classical computers capable of performing VQAs. The initial part of the algorithm distributes $R$ to Rob, $A$ to Alice, and $B$ to Bob, who are all in distant locations. Then, Rob performs the parameterized measurement and sends the outcome over a classical channel to Alice, who performs another parameterized measurement. Then, they can repeat this process to assess the quality of the entanglement between Alice and Bob. This interpretation is even more interesting regarding quantum networks for the multipartite case, in which the classical data gets broadcast from Rob to all the other nodes except the last one.
@@
VQSAs can tackle other problems involving quantum steering, like maximizing the pure-state decompositions of quantum states. This technique may also be helpful for estimating other entanglement measures that involve optimizing the set of separable states. By applying the insights of [61, Appendix A] and our approach here, it is clear that VQSAs will also be helpful for estimating maximal fidelities associated with other resource theories, such as the resource theory of coherence [62]. More broadly, we suspect that the paradigm of parameterized mid-circuit measurements and distributed variational quantum algorithms will be helpful in addressing other computational problems of interest in quantum information science and physics, given recent advances in experimental implementations [63, 64, 65, 66].
@@ 4 Methods
This section briefly overviews the techniques used to prove Theorem 1, our main result, a brief description of SDP benchmarks, and essential details about our simulations.
@@
Our pure-state separability test consists of a distributed quantum computation involving a prover and a verifier (see Figure 8). The computation starts with the verifier preparing the pure state $\psi_{AB}$. The prover sends the verifier the pure state $\phi_{A^{\prime}}$ in register $A^{\prime}$. (We note that the prover can send a mixed state; however, the maximum acceptance probability of the test is achieved by a pure state. Hence, without loss of generality, the prover must send a pure state.) The verifier then performs the standard swap test [70, 71] on $A$ and $A^{\prime}$ and accepts if the measurement outcome is zero. In the standard model of quantum computational complexity [41, 42], the prover attempts to get the verifier to accept the swap test with as high a probability as possible. Thus, in this scenario, the prover selects $\phi_{A^{\prime}}$ to maximize the overlap between the reduced stated $\psi_{A}\coloneqq\operatorname{Tr}_{B}[\psi_{AB}]$ and $\phi_{A^{\prime}}$. The maximum acceptance probability is then equal to
@@
Now we outline the proof of Theorem 1, which relies on two important facts. The first is that the fidelity of separability can be written in terms of a convex roof as follows [61, Theorem 1]:
@@
where $\{p(x)\}_{x}$ is a probability distribution and each $\psi_{AB}^{x}$ is a pure state. See also [72, Lemma 1]. The second fact is that, for a pure bipartite state, $F_{s}(\psi_{AB})$ can be rewritten as [61, Section 6.2]
@@
First, let us recall that the fidelity between two quantum states has an SDP formulation [73]. Since there is no semidefinite constraint that directly corresponds to optimizing over the set of separable states [74], we can approximate the fidelity of separability of a state by maximizing its fidelity with positive partial transpose (PPT) states [27, 28] and $k$-extendible states [29, 30]. Further noting that the PPT and $k$-extendibility constraints are positive semidefinite constraints, we obtain our first benchmark $\widetilde{F}_{s}^{1}(\rho_{AB},k)$, defined in Appendix E, and which is proven there to satisfy the following bounds:
@@
An essential issue with variational quantum techniques, such as VQAs, is the emergence of barren plateaus or vanishing gradients as the number of qubits increases [43]. However, recent results have shown that this problem can be mitigated by switching from a global reward function to a local reward function [44]. In our case, a global reward function is one in which we measure all the qubits that constitute system $A$, as done in the approach discussed in Theorem 2. An example of a local reward function involves selecting a qubit in the system $A$ at random to measure in the computational basis and recording the outcome, accepting if the result is equal to zero. Our proposed local reward function can be used to obtain upper and lower bounds on our initial global reward function, following the approach of [81, Appendix C] and discussed for completeness in Appendix L. Local functions have also been used to avoid barren plateaus in VQAs to determine the geometric measure of entanglement for pure states [82]. We provide simulations of the local reward function in Appendix G, indicating that the local reward function can also be used to estimate the fidelity of separability of a given state.

### v4

@@ 1 Introduction
where $\{|x\rangle_{R}\}_{x\in\mathcal{X}}$ is an orthonormal basis. Purifications are not unique, but all other purifications of $\sigma_{AB}$ are related to the one in (2) by the action of a unitary operation on the reference system $R$ [40]. By inspecting (2), we see that the systems $A$ and $B$ can be steered into the probabilistic ensemble $\{(p(x),\psi^{x}_{A}\otimes\phi^{x}_{B})\}_{x\in\mathcal{X}}$ of product states by performing the projective measurement $\{|x\rangle\!\langle x|_{R}\}_{x\in\mathcal{X}}$ on the reference system $R$ of $\varphi_{RAB}$. This leads to an idea for testing separability in the general case. If a purification of a general state $\rho_{AB}$ is available and the state $\rho_{AB}$ is indeed separable, then one can a) try to find the unitary that realizes the purification in (2) and b) perform the measurement $\{|x\rangle\!\langle x|_{R}\}_{x\in\mathcal{X}}$ on the reference system $R$. After receiving the outcome $x$, one can finally test whether the reduced state is a product state.
@@
In this paper, we expand on the abovementioned idea to develop various separability tests using the quantum steering effect. Our separability test for mixed states consists of a distributed quantum computation involving two parties: a computationally unbounded server, called a prover, which can, in principle, perform any quantum computation imaginable, and a computationally limited client, called a verifier, which can perform time-efficient quantum computations (see Figure 1). We prove Theorems 1 and 2, which state that the acceptance probabilities of our algorithms, in the ideal case, are directly related to a bonafide entanglement measure, the fidelity of separability. We also employ concepts from quantum computational complexity theory [41, 42] to understand how difficult this test is to perform. Our second contribution results from a modification of our separability test. In an attempt to design a practical algorithm, we replace the prover with a combination of parameterized unitary circuits and classical optimization techniques to perform the necessary computation. This results in a variational quantum steering algorithm (VQSA) that approximates the aforementioned separability test (see Figure 2). The concept of quantum steering is again at the heart of our VQSA, just like the test for separability that it approximates. Interestingly, we prove that the acceptance probability of both tests is related to an entanglement measure called fidelity of separability [22, 23]. We also generalize our separability test and VQSA to the multipartite setting using appropriate definitions of multipartite separability.
@@
Our approach is distinct from recent work on quantum algorithms for estimating entanglement. For example, VQAs have been used to address this problem by estimating the Hilbert–Schmidt distance [45], by creating a zero-sum game using parameterized unitary circuits [46], by employing symmetric extendibility tests [31], by estimating logarithmic negativity [47], and using the positive map criterion [47]. VQAs have also been used to estimate the geometric measure of entanglement of multiqubit pure states [48]. The work of [49] is the closest related to ours, but the test used there requires two copies of the state of interest and controlled swap operations for each run of the algorithm, while our VQSA does not require either. In contrast, we introduce a paradigm for VQAs involving parameterized mid-circuit measurements, which is the core of our method for estimating entanglement, and we suspect that this approach will be helpful in future work for a wide variety of VQAs. Furthermore, as we show in Theorems 1 and 2, the acceptance probabilities of our algorithms, in the ideal case, are directly related to a bonafide entanglement measure, the fidelity of separability.
@@ 2.1 Quantum Interactive Proof for Fidelity of Separability
where $\Pi_{A^{\prime}A}^{\operatorname{sym}}$ is the projector onto the symmetric subspace of the $A^{\prime}$ and $A$ systems, and $\operatorname{EB}_{R\to A^{\prime}}$ denotes the set of all entanglement-breaking channels with input system $R$ and output system $A^{\prime}$. We state in Theorem 1 below that the maximum acceptance probability in (4) can be expressed as a simple function of the fidelity of separability of $\rho_{AB}$, the latter defined as [22, 23]
@@
For a pure state $\psi_{RAB}$, the following equality holds:
@@ 2.2 Variational Quantum Steering Algorithm for Fidelity of Separability
Our VQSA is structured as follows. Let $\rho_{AB}$ denote the state for which we want to estimate the fidelity of separability, and let $\psi_{RAB}$ be a purification of it, which results from the action of the unitary operator $U^{\rho}$ on the all-zeros pure state $|0\rangle\!\langle 0|$. Once we have $\psi_{RAB}$, we can attempt to access all possible pure-state decompositions $\{(p(x),\psi_{AB}^{x})\}_{x\in\mathcal{X}}$ of $\rho_{AB}$ by acting on system $R$ with unitary operations. We use the first parameterized unitary $W_{R}(\Theta)$. To ensure that we have a sufficient number of measurement outcomes (to cover the possible case when $|\mathcal{X}|=\text{rank}(\rho_{AB})^{2}$), we can prepare some ancilla qubits in the all-zeros state of a system $R^{\prime}$ and act with $W$ on $R$ and $R^{\prime}$. However, without loss of generality, these extra qubits can be grouped as part of an overall reference system, relabeled as $R$.
@@ 2.3 Benchmarking Semidefinite Programs and Examples
As a second example, we consider a state consisting of four qubits. Let us consider the four-qubit state $|\psi\rangle$ defined as follows:
@@ 2.4 Generalization to Multipartite Fidelity of Separability
We also generalize our VQSA to measure the fidelity of separability of multipartite states in the following fashion.
A multipartite state $\rho_{A_{1}\cdots A_{M}}\in\mathcal{D}(\mathcal{H}_{A_{1}\cdots A_{M}})\equiv\mathcal{D}(\mathcal{H}_{A_{1}}\otimes\cdots\otimes\mathcal{H}_{A_{M}})$ is separable if it can be written as
@@
where $\{\mu^{x}_{R}\}_{x}$ is a rank-one POVM and $\{\phi^{x,i}_{A^{\prime}_{i}}\}_{x,i}$ is a set of pure states. The prover sends systems $(A^{M-1})^{\prime}\equiv A^{\prime}_{1}\cdots A^{\prime}_{M-1}$ to the verifier. (Here again, we can think of the prover as actually being split into $M$ provers, a first who performs the measurement $\{\mu^{x}_{R}\}_{x}$ and communicates the outcome $x$ to $M-1$ other provers, the $i$th of whom prepares the state $\phi^{x,i}_{A^{\prime}_{i}}$ and sends it to the verifier, for all $i\in\{1,\ldots,M-1\}$.) Finally, the verifier performs a collective swap test on these systems and the systems $A_{1}\cdots A_{M}$, as depicted in Figure 5. The acceptance probability of this distributed quantum computation is given by
@@
where $\Pi_{(A^{M-1})^{\prime}A^{M-1}}^{\operatorname{sym}}$ is the projection onto the symmetric subspace of systems $(A^{M-1})^{\prime}$ and $A^{M-1}$ and $\operatorname{EB}_{M-1}$ denotes the set of multipartite entanglement-breaking channels defined in (11). This leads to the following theorem, which generalizes Theorem 1 to the multipartite case:
@@
Figure 6: VQSA to estimate the multipartite fidelity of separability $F_{s}(\rho_{A_{1}A_{2}A_{3}A_{4}})$. The unitary circuit $U^{\rho}$ prepares the state $\psi_{RA_{1}A_{2}A_{3}A_{4}}$, which is a purification of $\rho_{A_{1}A_{2}A_{3}A_{4}}$. The parameterized circuit $W_{R}(\Theta)$ acts on $R$ to evolve the state to another purification of $\rho_{A_{1}A_{2}A_{3}A_{4}}$. The following measurement, labeled “steering measurement,” steers the remaining systems to be in a state $\psi_{A_{1}A_{2}A_{3}A_{4}}^{x}$ if the measurement outcome $x$ occurs. Conditioned on the outcome $x$, the final parameterized circuits $U^{x,1}_{A_{1}}(\Theta^{x}_{1})$, $U^{x,2}_{A_{2}}(\Theta^{x}_{2})$, and $U^{x,3}_{A_{3}}(\Theta^{x}_{3})$ are applied, and the subsequent measurement accepts with a maximum probability of $F_{s}(\rho_{A_{1}A_{2}A_{3}A_{4}})$.
@@ 2.5 Quantum Computational Complexity Considerations
$\max_{\begin{subarray}{c}\{(p(x),\psi^{x})\}_{x},\left\{\varphi^{x}\right\}_{x},\\ \rho_{B}=\sum_{x}p(x)\psi_{B}^{x}\end{subarray}}\sum_{x}p(x)F(\psi_{B}^{x},\mathcal{N}_{A\rightarrow B}(\varphi_{A}^{x})),$ (15)
where $\{(p(x),\psi^{x})\}_{x}$ is a pure-state ensemble and $\left\{\varphi^{x}\right\}_{x}$ is a set of pure states. See Appendix J for details and an interpretation of this quantity.
@@ 3 Conclusion and Discussion
In this paper, we detailed a distributed quantum computation to test the separability of a quantum state that, at its core, uses quantum steering. This test demonstrated a link between quantum steering and the separability problem. The acceptance probability of this distributed quantum computation is directly related to an entanglement measure known as the fidelity of separability. Using the test’s structure, we also showed computational complexity-theoretic results and established a link between quantum steering, quantum algorithms, and quantum computational complexity. By replacing the prover with a parameterized circuit, we modified this distributed quantum computation to develop our variational quantum steering algorithm (VQSA), a novel kind of variational quantum algorithm that uses quantum steering to address the problem of estimating the fidelity of separability. This algorithm allows for the direct estimation of the fidelity of separability without the need for state tomography and subsequent approximate tests on separability. Our algorithm is not unitary due to the mid-circuit measurement on system $R$ and the consequent conditional operation applied on system $A$. This is an important distinction from most VQAs, which do not use a parameterized mid-circuit measurement. We also discuss multipartite generalizations of both our separability test and VQSA. Finally, we simulated our VQSA using the noisy Qiskit Aer simulator [58], which showed favorable convergence trends and was compared against two classical SDP benchmarks.
@@
Our VQSA has applications beyond entanglement quantification on a single quantum computer. We can also think of our VQSA as a distributed variational quantum algorithm for measuring the entanglement of a bipartite state. See [59, 60, 61] for previous instances of distributed VQAs. Indeed, our algorithm can be executed over a quantum network in which each node has quantum and classical computers capable of performing VQAs. The initial part of the algorithm distributes $R$ to Rob, $A$ to Alice, and $B$ to Bob, who are all in distant locations. Then, Rob performs the parameterized measurement and sends the outcome over a classical channel to Alice, who performs another parameterized measurement. They can repeat this process to assess the quality of the entanglement between Alice and Bob. This interpretation is even more interesting regarding quantum networks for the multipartite case, in which the classical data gets broadcast from Rob to all the other nodes except the last one.
@@
VQSAs can tackle other problems involving quantum steering, like maximizing the pure-state decompositions of quantum states. This technique may also be helpful in estimating other entanglement measures that involve optimizing over the set of separable states. By applying the insights of [62, Appendix A] and our approach here, it is clear that VQSAs will also help estimate maximal fidelities associated with other resource theories, such as the resource theory of coherence [63]. More broadly, we suspect that the paradigm of parameterized mid-circuit measurements and distributed variational quantum algorithms will help address other computational problems of interest in quantum information science and physics, given recent advances in experimental implementations [64, 65, 66, 67].
@@ 4 Methods
This section briefly overviews the techniques used to prove Theorem 1 (one of our main results), a brief description of our SDP benchmarks, and essential details about our simulations.
@@
Our pure-state separability test consists of a distributed quantum computation involving a prover and a verifier (see Figure 8). The computation starts with the verifier preparing the pure state $\psi_{AB}$. The prover sends the verifier the pure state $\phi_{A^{\prime}}$ in register $A^{\prime}$. (We note that the prover can send a mixed state; however, the maximum acceptance probability of the test is achieved by a pure state. Hence, without loss of generality, the prover should send a pure state.) The verifier then performs the standard swap test [71, 72] on $A$ and $A^{\prime}$ and accepts if the measurement outcome is zero. In the standard model of quantum computational complexity [41, 42], the prover attempts to get the verifier to accept the swap test with as high a probability as possible. Thus, in this scenario, the prover selects $\phi_{A^{\prime}}$ to maximize the overlap between the reduced stated $\psi_{A}\coloneqq\operatorname{Tr}_{B}[\psi_{AB}]$ and $\phi_{A^{\prime}}$. The maximum acceptance probability is then equal to
@@
Now we outline the proof of Theorem 1, which relies on two important facts. The first is that the fidelity of separability can be written in terms of a convex roof as follows [62, Theorem 1]:
@@
where $\{p(x)\}_{x}$ is a probability distribution and each $\psi_{AB}^{x}$ is a pure state. See also [73, Lemma 1]. The second fact is that, for a pure bipartite state, $F_{s}(\psi_{AB})$ can be rewritten as [62, Section 6.2]
@@
First, let us recall that the fidelity between two quantum states has an SDP formulation [74]. Since there is no semidefinite constraint that directly corresponds to optimizing over the set of separable states [75], we can approximate the fidelity of separability of a state by maximizing its fidelity with positive partial transpose (PPT) states [27, 28] and $k$-extendible states [29, 30]. Further noting that the PPT and $k$-extendibility constraints are semidefinite constraints, we obtain our first benchmark $\widetilde{F}_{s}^{1}(\rho_{AB},k)$, defined in Appendix E, and which is proven there to satisfy the following bounds:
@@
An essential issue with variational quantum techniques, such as VQAs, is the emergence of barren plateaus or vanishing gradients as the number of qubits increases [43]. However, recent results have shown that this problem can be mitigated by switching from a global reward function to a local reward function [44]. In our case, a global reward function is one in which we measure all the qubits that constitute system $A$, as done in the approach discussed in Theorem 2. An example of a local reward function involves selecting a qubit in the system $A$ at random to measure in the computational basis and recording the outcome, accepting if the result equals zero. Our proposed local reward function can be used to obtain upper and lower bounds on our initial global reward function, following the approach of [82, Appendix C] and discussed for completeness in Appendix L. Local functions have also been used recently to avoid barren plateaus in VQAs when estimating the geometric measure of entanglement for pure states [83]. We provide simulations of the local reward function in Appendix G, indicating that the local reward function can also be used to estimate the fidelity of separability of a given state.